# Sparsity May Be All You Need: Sparse Random Parameter Adaptation

### v1

@@ Appendix B Additional Dataset details
Dataset Classes Train Dev Test
COLA 2 7,551 1,000 1,043
MNLI 3 100,000 10,000 19,647
MRPC 2 3,668 408 1,725
QNLI 2 99,701 5,000 5,463
QQP 2 100,000 5,000 40,430
RTE 2 2,165 300 277
SST-2 2 66,349 1,000 872
IMDB 2 25,000 5,000 20,000
Table 9: Sequence classification datasets. Training sets limited to 100K samples. Training samples with more than 256 tokens are removed (here using mistral-7b tokenizer).
IMDB: contains a sample of “highly polar” movie reviews obtained from the online Internet Movie Database (IMDb) website. IMDb registered users provide a rating (from $1$ to $10$) with each review. The reviews are binary (positive/negative) labeled with their sentiment, defined from user ratings (Maas et al., 2011). Reviews with a rating higher or equal than $7$ are given a positive label; and a negative label if the rating is lower or equal than $4$. No reviews with ratings beyond these ranges are present in the dataset, which was constructed to have an equal number of positive and negative reviews, so guessing randomly yields 50% accuracy.
GLUE: Refer to Wang et al. (2019) which contains descriptions and details on all GLUE datasets.
Table 9 shows the actual splits used in our experiments for the IMDB and GLUE datasets, where the training data was filtered using the mistral-7b tokenizer by a threshold in the maximum token length of an network input. The only differences are for the training splits of the MRPC, QNLI, and RTE datasets.
@@ Appendix C Training details
We tokenized all training examples before starting the fine-tuning and looked at the histogram of their token lengths. To avoid batches with too much padding and improve training efficiency, we dropped those training examples with a disproportionate large token length, i.e., corresponding to a tail of extreme values in the histogram. We only do this for the training data since evaluation (on the development or test data) requires much less compute and memory (no gradients need to be calculated and stored) and is performed less frequently. We will indicate in Tables 2 and 9 the final splits after filtering the training data this way, with the final number of training examples used for the fine-tuning.
@@ Appendix D Additional Experiment Results
Tables 10 and Table 11 present the complete set of results on GLUE for the PT base models gemma-2b and mistral-7b, respectively.
@@
Model: google/gemma-2-2b QNLI RTE SST2 QQP MNLI MRPC COLA loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc Full FT 0.20 92.6 0.43 80.1 0.11 96.7 0.28 88.3 0.36 86.4 0.33 86.1 0.41 60.3 SpaRTA (5%) 0.15 94.4 0.46 80.5 0.11 96.8 0.23 90.1 0.28 89.4 0.38 83.4 0.34 64.6 SpaRTA (0.5%) 0.15 94.1 0.61 69.0 0.12 96.3 0.24 89.8 0.30 89.0 0.38 83.4 0.38 62.2 SpaRTA (0.037%) 0.29 88.3 0.69 57.8 0.11 96.2 0.30 86.9 0.47 81.1 0.58 70.3 0.52 42.1 LoRA (0.037%) 0.17 93.6 0.59 69.0 0.11 96.1 0.26 88.6 0.30 88.6 0.36 84.5 0.36 61.7 Head (1.63e-4%) 0.56 71.8 0.68 60.3 0.32 89.1 0.43 79.9 0.89 58.7 0.59 70.4 0.53 33.2
@@
Table 10: Test loss and accuracy of model google/gemma-2-2b adapted to glue datasets with different fine-tuning methods. For training details see Table 13.
@@
Model: mistralai/Mistral-7B-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc Full FT 0.12 95.4 0.28 89.9 0.11 96.9 0.23 90.1 0.25 91.1 0.34 85.9 0.32 66.8 SpaRTA (5%) 0.12 95.5 0.34 86.3 0.11 96.9 0.23 90.3 0.25 90.9 0.34 87.6 0.37 64.3 SpaRTA (0.5%) 0.12 95.9 0.35 86.3 0.11 96.7 0.22 90.5 0.24 91.2 0.32 88.8 0.34 68.2 SpaRTA (0.048%) 0.13 95.3 0.36 86.3 0.11 96.8 0.24 89.7 0.26 90.3 0.38 86.3 0.34 66.9 LoRA (0.048%) 0.12 95.6 0.30 88.1 0.11 96.7 0.23 90.0 0.25 91.0 0.31 87.3 0.32 67.8 Head (1.15e-4%) 0.39 82.8 0.63 70.4 0.16 94.8 0.33 85.8 0.63 74.0 0.53 73.8 0.48 45.5
@@
Table 11: Test loss and accuracy of model mistralai/Mistral-7B-v0.3 adapted to glue datasets with different fine-tuning methods. For training details see Table 13.
@@ E.1 IMDB
E.2 GLUE
For GLUE, given the number of tasks (7) and the number of methods tried (6), reporting each best set of training parameters would be just prohibitive. Instead, the trainings of these methods was done using a simple grid search over some of the hyper-parameters. The range on these hyper-parameters are reported in Table 13. Similarly to IMDB, any sample longer than 256 tokens (which is dependnent on the tokenizer used) is discarded from the dataset to avoid having few minibatch with one very long sample compared to the rest, as seen in Table 2.
Parameter Grid Search Value Range
Full FT batch size [32, 24, 8] depending on task and model to fit in GPU memory
learning rate [1e-3, 5e-4, 2e-4, 1e-4, 5e-5, 1e-5, 5e-6]
dropout [0.1]
weight decay [0.1]
SpaRTA batch size [same as for Full FT]
learning rate [same as for Full FT]
weight decay [0.1]
LoRA batch size [same as for Full FT]
learning rate [same as for Full FT]
$r$ [8]
$\alpha$ [16]
dropout [0.1]
weight decay [0.1]
Head batch size [same as for Full FT]
learning rate [same as for Full FT]
dropout [0.1]
weight decay [0.0]
Table 13: (GLUE) Training parameters and their grid search value range to get results for each adaptation method for each model in Table 4, Table 5, Table 10, and Table 11. Best set of parameters were chosen from best validation loss results over 2 epochs of training.

### v2

@@ Appendix B Additional Dataset details
IMDB contains a sample of “highly polar” movie reviews obtained from the online Internet Movie Database (IMDb) website. IMDb registered users provide a rating (from $1$ to $10$) with each review. The reviews are binary (positive/negative) labeled with their sentiment, defined from user ratings (Maas et al., 2011). Reviews with a rating higher or equal than $7$ are given a positive label; and a negative label if the rating is lower or equal than $4$. No reviews with ratings beyond these ranges are present in the dataset, which was constructed to have an equal number of positive and negative reviews, so guessing randomly yields 50% accuracy.
GLUE datasets are described in details in Wang et al. (2019).
BoolQ Clark et al. (2019) is a reading comprehension dataset with binary yes/no questions. Each example is a triplet of (passage, question, answer). It represents a natural language inference task in which a passage and a question are given as input, and a yes or no answer should be predicted.
MMLU Hendrycks et al. (2020) is a benchmark designed to measure the knowledge a model acquires during pre-training. To facilitate the elicitation of a model’s knowledge using multiple-choice questions, MMLU also comes with a dataset of examples on multiple-choice question answering. Thus, this training set is not designed to increase the model knowledge about the world, but to teach a model to answer knowledge questions in multiple-choice format in a zero-shot setting.
@@ Appendix C Training details
We tokenized all training examples before starting the fine-tuning and looked at the histogram of their token lengths. To avoid batches with too much padding and improve training efficiency, we dropped those training examples with a disproportionate large token length, i.e., corresponding to a tail of extreme values in the histogram. We only do this for the training data since evaluation (on the development or test data) requires much less compute and memory (no gradients need to be calculated and stored) and is performed less frequently. We indicate in Tables 2 the final splits after filtering the training data this way, with the final number of training examples used for the fine-tuning.
@@ Appendix D Additional Experiment Results
Tables 10 and Table 11 present the complete set of results on GLUE and BoolQ for the PT base models gemma-2b and mistral-7b, respectively. Tables 12 and 13 present standard errors from corresponding experiments.
@@
Model: gemma-2b QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ Method Targets loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc loss acc. SpaRTA (5%) ALL 0.15 94.5 0.50 80.5 0.12 96.6 0.23 90.2 0.39 85.5 0.38 84.6 0.36 62.6 0.32 86.7 SpaRTA (0.5%) 0.15 94.5 0.51 77.3 0.12 96.1 0.23 90.1 0.42 84.1 0.37 84.3 0.36 62.3 0.34 85.4 SpaRTA (0.037%) 0.20 91.9 0.62 66.3 0.13 95.4 0.27 88.3 1.34 49.3 0.51 76.1 0.49 50.7 0.39 82.8 SpaRTA (0.037%) Q,V 0.18 93.0 0.58 70.4 0.13 95.7 0.26 88.8 1.04 56.8 0.43 80.3 0.41 56.9 0.36 84.5 LoRA (0.037%) 0.15 94.2 0.62 65.5 0.11 96.8 0.26 88.9 0.42 84.0 0.47 77.3 0.40 60.9 0.34 85.7 DoRA (0.037%) 0.15 94.4 0.62 66.5 0.11 96.6 0.25 89.0 0.41 84.4 0.47 77.6 0.39 60.6 0.43 78.9 SpaRTA (0.037%) O,V 0.16 93.8 0.55 73.2 0.11 96.5 0.25 89.3 0.51 80.4 0.39 84.3 0.39 61.9 0.34 86.1
@@
Table 10: Test loss and accuracy of model gemma-2b adapted to GLUE and BoolQ datasets with different fine-tuning methods. Results are averaged over 3 random seeds. For training details see Appendix E.2.
@@
Model: Mistral-7B-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method Targets loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc loss acc. SpaRTA (5%) ALL 0.11 95.9 0.36 85.0 0.11 96.8 0.38 82.1 0.30 89.2 0.33 86.7 0.35 64.6 0.25 90.1 SpaRTA (0.5%) 0.12 95.8 0.44 84.4 0.11 96.7 0.22 90.6 0.26 90.4 0.32 87.0 0.35 66.8 0.26 89.9 SpaRTA (0.048%) 0.13 95.3 0.41 83.6 0.11 96.7 0.23 90.1 0.40 86.7 0.34 86.6 0.36 64.6 0.26 89.8 SpaRTA (0.048%) Q,V 0.14 94.5 0.37 84.0 0.11 96.6 0.24 89.7 0.33 88.6 0.36 85.5 0.36 65.2 0.28 89.3 LoRA (0.048%) 0.12 95.4 0.39 85.4 0.12 96.5 0.23 90.1 0.28 89.7 0.32 87.0 0.38 64.5 0.27 89.6 DoRA (0.048%) 0.12 95.6 0.41 85.3 0.12 96.8 0.24 90.1 0.28 89.7 0.33 86.4 0.34 65.8 0.26 89.9 SpaRTA (0.048%) O,V 0.15 94.5 0.41 85.6 0.21 91.7 0.24 89.6 0.33 87.8 0.35 86.4 0.39 59.8 0.29 89.3
@@
Table 11: Test loss and accuracy of model mistral-7b adapted to GLUE and BoolQ datasets with different fine-tuning methods. Results are averaged over 3 random seeds. For training details see Appendix E.2.
Model: gemma-2b-it QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc loss acc. PT (zero-shot) 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 SpaRTA (5%) ALL 0.00 0.14 0.03 1.54 0.01 0.51 0.00 0.21 0.01 0.26 0.01 0.48 0.00 0.88 0.01 0.54 SpaRTA (0.5%) 0.00 0.16 0.03 0.64 0.01 0.46 0.00 0.17 0.00 0.13 0.01 0.07 0.01 0.75 0.01 0.25 SpaRTA (0.037%) 0.00 0.12 0.01 0.55 0.00 0.17 0.00 0.18 0.00 0.14 0.00 0.45 0.01 0.44 0.00 0.09 LoRA (0.037%) Q,V 0.00 0.11 0.01 1.10 0.00 0.35 0.00 0.03 0.00 0.11 0.00 0.22 0.02 1.40 0.01 0.12 DoRA (0.037%) 0.00 0.11 0.01 0.55 0.00 0.20 0.00 0.03 0.00 0.09 0.00 0.32 0.01 1.31 0.01 0.29 SpaRTA (0.037%) 0.00 0.08 0.00 1.04 0.01 0.11 0.00 0.00 0.00 0.26 0.00 0.22 0.02 2.83 0.00 0.41 SpaRTA (0.037%) O,V 0.00 0.13 0.01 0.64 0.00 0.27 0.00 0.16 0.00 0.17 0.02 0.22 0.01 4.36 0.00 0.22
Table 12: Standard errors of test loss and accuracy of model gemma-2b-it adapted to GLUE and BoolQ datasets with different fine-tuning methods.
Model: Mistral-7B-Instruct-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc loss acc. PT (zero-shot) 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 SpaRTA (5%) ALL 0.00 0.09 0.01 0.24 0.01 0.26 0.00 0.03 0.00 0.08 0.01 0.47 0.01 1.91 0.00 0.16 SpaRTA (0.5%) 0.00 0.05 0.01 1.07 0.01 0.18 0.00 0.10 0.00 0.09 0.02 0.25 0.01 1.28 0.00 0.09 SpaRTA (0.048%) 0.00 0.10 0.03 0.55 0.00 0.20 0.00 0.03 0.00 0.11 0.01 0.32 0.02 1.87 0.00 0.20 LoRA (0.048%) Q,V 0.00 0.16 0.01 0.60 0.01 0.44 0.00 0.06 0.00 0.13 0.01 0.33 0.01 1.16 0.00 0.17 DoRA (0.048%) 0.02 0.76 0.02 0.87 0.00 0.17 0.00 0.03 0.00 0.07 0.00 0.17 0.01 1.22 0.00 0.16 SpaRTA (0.048%) 0.00 0.06 0.01 0.73 0.00 0.23 0.00 0.04 0.00 0.05 0.01 0.43 0.01 1.17 0.00 0.27 SpaRTA (0.048%) O,V 0.00 0.02 0.01 0.42 0.01 0.34 0.00 0.07 0.00 0.13 0.02 0.24 0.01 0.97 0.00 0.23
Table 13: Standard errors of test loss and accuracy of model mistral-7b-it adapted to BLUE and BoolQ datasets with different fine-tuning methods.
@@ E.1 IMDB
Model: google/gemma-2b-it method QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ SpaRTA (5%) 1e-5 5e-4 5e-5 5e-5 1e-5 2e-4 5e-5 5e-5 SpaRTA (0.5%) 2e-4 1e-3 2e-4 2e-4 2e-4 5e-4 1e-3 2e-4 SpaRTA (0.037%) 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 LoRA (0.037%) 2e-4 1e-3 5e-4 5e-4 5e-4 2e-4 5e-4 5e-4 DoRA (0.037%) 2e-4 1e-3 5e-4 5e-4 5e-4 2e-4 5e-4 5e-4 SpaRTA (0.037%) 5e-4 1e-3 5e-4 5e-4 5e-4 5e-4 1e-3 5e-4 SpaRTA (0.037%) 5e-4 1e-3 5e-4 5e-4 5e-4 5e-4 1e-3 5e-4
Table 14: Optimal learning rates for gemma-2b-it model for different fine-tuning methods.
Model: Mistral-7B-Instruct-v0.3 method QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ SpaRTA (5%) 1e-5 5e-6 5e-6 1e-5 1e-5 5e-5 5e-5 5e-6 SpaRTA (0.5%) 2e-4 1e-4 5e-5 5e-4 2e-4 5e-4 2e-4 5e-5 SpaRTA (0.048%) 5e-4 5e-4 5e-4 1e-3 5e-4 1e-3 1e-3 1e-3 LoRA (0.048%) 2e-4 1e-4 2e-4 2e-4 1e-4 2e-4 2e-4 1e-4 DoRA (0.048%) 2e-4 1e-4 2e-4 2e-4 1e-4 2e-4 2e-4 1e-4 SpaRTA (0.048%) 2e-4 1e-4 2e-4 5e-4 2e-4 1e-3 1e-3 5e-4 SpaRTA (0.048%) 2e-4 1e-4 2e-4 5e-4 2e-4 1e-3 1e-3 5e-4
Table 15: Optimal learning rates for mistral-7b-it model for different fine-tuning methods.
@@
E.2 GLUE and BoolQ
For GLUE and BoolQ, training was done using a simple grid search over the learning rates: [1e-3, 5e-4, 2e-4, 1e-4, 5e-5, 1e-5, 5e-6]. The optimal learning rates found are given in Tables 14 and 15. Weight decay was set to 0. Dropout was set to 0.0 for SpaRTA and 0.1 for LoRA and DoRA. Batch sizes were set according to what could fit in GPU memory (different per dataset and model). The number of epochs was set to 3 for instruct models and 2 for base models. Similarly to IMDB, any sample longer than 256 tokens (which is dependnent on the tokenizer used) is discarded from the dataset to avoid having few minibatch with one very long sample compared to the rest, as seen in Table 2.
E.3 MMLU
As with other datasets, we started approximating the deltas of the self-attention matrices Wq and Wv with rank $r=8$ matrices, resulting in approximately 1M trainable parameters for the Gemma 2B models and 3.5M for the Mistral 7B models under LoRA ($\alpha=16$). We choose the sparsity of our SpaRTA adapter accordingly so both methods end up with the same number of parameters to train. Thus, our SpaRTA approach uses a $s=99.96\%$ for the Gemma models (i.e. making only $0.04\%$ of the original parameters trainable); and $s=99.95\%$ for the Mistral models.
Since we observed that gemma-2b struggled to learn with both SpaRTA and LoRA methods, we decided to increase the number of trainable parameters when adapting gemma-2b to MMLU. Specifically, we increased the rank of the LoRA adaption matrices to $r=16$, which lead to approximately 2M trainable parameters; and chose accordingly a sparsity $s=99.92\%$ for SpaRTA.
The MMLU dataset has a small fraction of training examples that are extremely long. We enforce a maximum input token length of $520$ for training efficiency. This reduces the number of training examples from $99,842$ to $84,296$ and $74,100$ for the Gemma and Mistral instruction models; and to $91,321$ and $85,820$ for their respective base models. Test and validations sets are not affected by this decision. The training parameters used for adapting each model to MMLU are shown in Table 17.
Training parameter gemma-2b gemma-2b-it mistral-7b mistral-7b-it
SpaRTA  / LoRA epochs 14 / 6 1 1 1
batch size 40 40 40 40
learning rate 1e-4 1e-4 5e-5 5e-5
dropout 0.0 0.05 0.1 0.1
weight decay 0.0 0.0 0.0 0.0
Table 17: Training parameters used with both SpaRTA and LoRA for each of the pre-trained models in Table 7 (MMLU).